Karen needs to prepare povidone-iodine solution in the correct concentration to prepare the surgical site. The...

  1. Karen needs to prepare povidone-iodine solution in the correct concentration to prepare the surgical site. The solution comes in a 7.5% concentration. Karen needs to dilute the solution to a concentration of 1%. She needs 100ml of solution to prepare the site.
    1. Identify the standard formula for preparing solutions. (1 point)
    2. How many milliliters of the 7.5% povidone-iodine should Karen start with? (2 points)
    3. How many milliliters of water does Karen need? (2 points)
Answer #1

A) [Initial volume - V1] [Initial concentration- C1] = [Final volume- V2] [Final concentration- C2]

B) Final concentration = 1%
   Final volume = 100 ml

Initial concentration = 7.5%

Initial volume = ?

V1 = (V2 C2) / C1

V1 = (100 * 1) / 7.5 = 13.33 ml

13.33 ml of 7.5% povidone-iodine solution is needed.

C) The final volume needed is 100 ml of which povidone-iodine solution is 13.33 ml and remaining is water. So

100-13.33 = 86.67 ml

• 13.33 ml of 7.5% povidone-iodine solution and 86.67 ml water make 1% 100 ml povidone-iodine solution.